OpenStudy (anonymous):

can someone tell us how to simplify 4x +8-5x

OpenStudy (anonymous):

=8-x

OpenStudy (anonymous):

4x-5x = x so the whole thing = 8-x

OpenStudy (anonymous):

can someone explain how please

OpenStudy (anonymous):

you can only simplify like terms there are two terms in x : 4x - 5x : works out to -x

OpenStudy (anonymous):

9-8a+7.5-4a

OpenStudy (anonymous):

is this one 16.5-12a?

OpenStudy (anonymous):

can someone tell us if we're on the right track?

myininaya (myininaya):

thats right

OpenStudy (anonymous):

super thanks!